# Generalized Uncertainty Principle for Entangled States of Two Identical Particles

### v3

@@ Abstract
In this work we determine the consequences of the quantum entanglement of a system of two identical particles when the generalized uncertainty principle (GUP) is considered. GUP is usually associated with the existence of a minimal length. We focus on the main formulations of the GUP and then we determine the minimal uncertainties in position induced by those modified GUP’s. Our results point out that the minimal uncertainty is reduced by half of its usual value independently of the GUP employed. This implies that the minimal length is also reduced by half. On the other hand, it is generally expected that the minimal length must not depend on physical system. We overcome this apparent paradox by realizing that the entangled system is composed by two particles so that an effective parameter related to the minimal length must be employed.
@@ 1 Introduction
Amongst the (many) new concepts introduced by the quantum mechanics, the quantum entanglement [1, 2] is one that, probably, more contradicts our common sense. Although, in the beginning, the quantum entanglement were only associated to theoretical aspects of quantum mechanics, specially those related to the non-locality or the complementarity (hidden variables) [3], nowadays it is a key component of the applications and experiments on quantum information, quantum computation and quantum teleportation [4, 5].
@@
The uncertainty principle is one of the fundamental cornerstones of the quantum mechanics. Nevertheless, it is a principle: its correct form can not be proven. That opens the way to the possibility that its canonical form, described by the Heisenberg’s uncertainty principle (HUP), can be generalized [^1] For more information on generalization of the Heisenberg’s uncertainty principle see Refs. [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16].. An example of the possible generalizations of HUP, whose origin can be traced back to quantum gravity, is given by introducing a non-zero minimal uncertainty in the measurement of position. That non-zero minimal uncertainty in position is, then, understood as a minimal-length scale, below which the necessary amount of energy to probe the position of a particle is so hight that it disturbs the space-time so that the concept of a length measurement loses its meaning. Hence, theories searching to describe a quantum approach for gravity lead generally to the existence of a minimal length. In fact, a minimal length actually appears in almost all proposed theories of the quantum gravity. For this reason theories formulated in a minimal-length scenario are considered to be effective theories of quantum gravity [17, 18, 19, 20].
@@
In 1999, Yoon-Ho Kim and Yanhua Shih conducted an experiment whose results apparently suggested a violation of the HUP [21]. However, G. Rigolin pointed out that in fact there is no violation of the HUP, because the HUP is derived for particles non-correlated (non-entangled). In the Kim and Shih’s experiment, the photons of the pair are correlated (entangled) when one of the physical slits is replaced by a virtual slit [^2] The interaction of the photon with a physical slit destroys the correlation between the photons of the pair.: the canonical HUP is no longer applicable since the quantum entanglement modifies the canonical HUP [22, 23].
@@
An immediate question arising from the previous considerations is how the quantum entanglement modifies a generalized uncertainty principle (GUP), in others words, which is the effect of the quantum entanglement in the minimal-length scale. The answer for that question is important in order to know the role of the quantum entanglement at the minimal-length scale (maybe appearing in the Planck scale) or in the early Universe. Unfortunately, this issue has been little considered in the literature. In [24], G. Blado, F. Herrera and J. Erwin have studied the inseparability conditions with the most usual GUP correction, whereas D. Park has used a coupled harmonic oscillator in order to find the effects of the quantum entanglement with a linear GUP in [25].
@@
The outline of this paper is as follows. In Section 2 we obtain an expression of the uncertainty principle for entangled states which is independent of the chosen GUP. In Section 3 we find the modified uncertainty principle for a pair of identical particles regarding the main proposals of GUP’s: Kempf, Mangano and Mann GUP (KMM-GUP), Ali, Das and Vagenas GUP (ADV-GUP), Pedram GUP and exponential all orders GUP. In Section 4 we estimate an upper bound for the minimal-length value. We present our conclusions in the Section 5.
@@ 2 Uncertainty principle for entangled states
The position and momentum linear operators of the i-th particle which act on the state vectors $\lvert\psi\rangle\in\cal{E}$ are the extensions $\tilde{Q}_{i}$ and $\tilde{P}_{i}$ defined as
@@
The relation (6) is general. It does not depend whether the system of particle is entangled or not.
@@
We briefly review the Rigolin’s result for a two particles system. From the definitions of $\Delta Q$ and $\Delta P$ we have [^3] Note that $[\tilde{Q}_{1},\tilde{Q}_{2}]=0$ and $[\tilde{P}_{1},\tilde{P}_{2}]=0$.
@@
From now on we omit the subscript $\psi$ for the sake of simplicity, whenever this does not cause any confusion. Thus,
@@
We now use the functions
which are called quantum covariance functions (QCF). By definition, QCF’s vanish if and only if the system is separable [28]. Therefore (11) and (12) are zero for any not entangled quantum system.
Using the QCF’s (11) and (12) we have
@@ 3 Uncertainty principle for entangled states in different minimal-length scenarios
In this section we consider a system of two entangled identical particles whose momenta have the same value but opposite directions, that is, $\vec{p}_{1}=-\vec{p}_{2}$, just as in the Kim and Shih’s experiment [21]. Therefore, in this case $\left\langle\hat{p}_{1}\right\rangle+\left\langle\hat{p}_{2}\right\rangle=0$. Moreover, such a consideration also allows us to estimate, in the next section, an upper bound for the value of the minimal length based on the experimental results obtained by Kim and Shih.
@@ 3.1 Heisenberg uncertainty principle
Hence
@@
Substituting Eq. (23) into Eq. (20) we get
$\Delta Q_{i}\Delta P_{i}\geq\frac{\hbar}{4}.$ (24)
The result (24) shows that for a system of two entangled identical particles the HUP is modified. Such an outcome is not new, it was already obtained by G. Rigolin in 2002 [22] and then in 2016 [23].
@@ 3.2 KMM GUP
$[\tilde{Q},\tilde{P}]=i\hbar\left[1+\beta\left(\tilde{P}_{1}^{2}+\tilde{P}_{2}^{2}\right)\right]=i\hbar\left(1+2\beta\tilde{P}_{i}^{2}\right).$ (27)
@@ 3.3 ADV GUP
$[\tilde{Q},\tilde{P}]=2i\hbar\left[1-\alpha\left(\tilde{P}_{1}+\tilde{P}_{2}\right)+2\alpha^{2}\left(\tilde{P}_{1}^{2}+\tilde{P}_{2}^{2}\right)\right].$ (31)
@@
$\langle[\tilde{Q},\tilde{P}]\rangle=2i\hbar\left[1-\alpha\left(\left\langle\tilde{P}_{1}\right\rangle+\left\langle\tilde{P}_{2}\right\rangle\right)+2\alpha^{2}\left(\left\langle\tilde{P}_{1}^{2}\right\rangle+\left\langle\tilde{P}_{2}^{2}\right\rangle\right)\right],$ (32)
$\langle[\tilde{Q},\tilde{P}]\rangle=2i\hbar\left[1+4\alpha^{2}\left\langle\tilde{P}_{i}^{2}\right\rangle\right].$ (33)
@@
where $\gamma\prime:=4\alpha^{2}\left\langle\tilde{P}_{i}\right\rangle^{2}$.
@@
It is important to note that the linear GUP (ADV-GUP) becomes non-linear in this case and consequently a maximal uncertainty in the momentum is no longer induced.
@@ 3.4 Pedram GUP
In order to overcome some problems arising from KMM-GUP and ADV-GUP - such as incorporation of a maximal momentum required in doubly special relativity (DSR) theories, commutative geometry and an approach valid for all order in the parameter related to the minimal length - P. Pedram has proposed a GUP [31, 32] based on the commutation relation given by
@@
$\Delta Q_{i}\Delta P_{i}\geq\frac{\hbar}{4}\frac{1}{\left[1-\beta\left(\Delta P_{i}\right)^{2}-\gamma\right]}.$ (39)
@@
3.5 Exponential all orders GUP
In the canonical field theory in the context of non-commutative coherent states representation and field theory on non-anticommutative superspace the Feynman propagator displays an ultra-violet (UV) cut-off of the form $e^{-\beta p^{2}}$ [34, 35, 36, 37]. In consequence, K. Nouicer has proposed an exponential all orders GUP [38, 39] based on the commutation relation given by
@@ 3.5 Exponential all orders GUP
$\Delta Q_{i}\Delta P_{i}\geq\frac{\hbar}{4}e^{\beta\left(\Delta P_{i}\right)^{2}+\gamma}.$ (44)
Therefore, the modified exponential all order GUP (44) introduces a minimal uncertainty given by
@@
Finally, we note that $\left(\Delta Q_{i}\right)_{min}$ is also twice smaller than for a non-entangled system of a pair of identical particles, as we have already said.
@@ 3.6 Minimal length
If a non-zero minimal uncertainty in position can be interpreted as a minimal length then the previous results show that the minimal length for two entangled identical particles can be twice smaller than for a separable system. It is clear this statement must be taken with care because a minimal length should be a constant, that is, it must not depend on the physical system, it is a quantum gravitation effect. In fact, the minimal length should be an invariant as well as the light speed is. The answer for that apparent contradiction is possibly because the system is made up of two particles. In references [40], C. Quesne and V. M. Tkachuck have claimed that for a system composed by $N$ particles the effective parameter related to the minimal length, $\beta$, is reduced by a factor $\frac{1}{N^{2}}$. Hence, $\frac{\hbar}{2}\sqrt{\beta}$ is not the correct minimal uncertainty in position $\left(\Delta Q_{i}\right)_{min}$ for the modified KMM-GUP, but $\frac{\hbar}{2}\sqrt{\beta_{i}}$, where [40, 41]
@@
The reader might want to claim that the minimal length is actually described by the minimal uncertainty of the entangled system. However, according to reference [23] we can presume that the minimal uncertainty for a system of $N$ entangled identical particles is reduced by $\frac{1}{N}$. Since, in principle, there is no limit for the number of particles for a entangled system, there would also be no limit for the minimal length.
@@ 4 Upper bound for the minimal-length value
$0.16\Delta P^{s}_{2}=\frac{\hbar}{2}\left[1+4\beta\left(\Delta P^{s}_{i}\right)^{2}+\gamma\right].$ (48)
@@
[…]
@@ 5 Conclusion
In this work we find the non-zero minimal uncertainties induced by the main proposals of GUP’s (KMM, ADV, Pedram and exponential) which are modified due to the quantum entanglement of a system of two identical particles. In principle, our results have pointed out that the minimal uncertainties are reduced at half for a system of two entangled identical particles independently of the GUP. Hence, if a non-zero minimal uncertainty in position can be interpreted as a minimal length then the quantum entanglement reduces by half the minimal length. However, the minimal length must not depend on the physical system. We overcome this apparent paradox by using the Quesne and Tkachuck’s proposal for a system composed. Consequently, despite the quantum entanglement to change the GUP, the minimal length does not change. Based on our results and using the reference [23] we can expect that the apparent minimal uncertainty for a system of $N$ entangled identical particles is reduced by $\frac{1}{N}$, nonetheless the minimal length does not change because the effective parameter $\beta$ is also reduced by a factor $\frac{1}{N^{2}}$.

### v4

@@ Abstract
In this work we determine the consequences of the quantum entanglement of a system of two identical particles when a generalized uncertainty principle (GUP) is considered. GUP’s are usually associated with the existence of a minimal length. We focus on the main GUP’s (KMM, ADV, Pedram and Nouicer) and then we determine the minimal uncertainties in position induced by those modified GUP’s. Our results point out that the minimal uncertainty is reduced by half of its usual value independently of the GUP employed. This implies that the minimal length is also reduced by half. On the other hand, it is generally expected that the minimal length must not depend on physical system. We overcome this apparent paradox by realizing that the entangled system is composed by two particles so that an effective parameter related to the minimal length must be employed.
@@ 1 Introduction
Amongst the (many) new concepts introduced by the quantum mechanics, the quantum entanglement [1, 2] is one that, probably, more contradicts our common sense. Although in the beginning the quantum entanglement were only associated to theoretical aspects of the quantum mechanics, specially those related to the non-locality or the complementarity (hidden variables) [3], nowadays it is a key component of the applications and experiments on quantum information, quantum computation and quantum teleportation [4, 5].
@@
The uncertainty principle is one of the fundamental cornerstones of the quantum mechanics. Nevertheless, it is a principle: its correct form can not be proven. This opens the way to the possibility that its canonical form, described by the Heisenberg’s uncertainty principle (HUP), can be generalized [^1] For more information on generalization of the Heisenberg’s uncertainty principle see Refs. [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16].. Possible generalizations of the HUP, whose origin can be traced back to quantum gravity, are given by introducing a non-zero minimal uncertainty in the measurement of position. The non-zero minimal uncertainty in position is, therefore, understood as a minimal-length scale, below which the necessary amount of energy to probe the position of a particle is so hight that it disturbs the space-time so that the concept of a length measurement loses its meaning. Hence, theories searching to describe a quantum approach for gravity lead generally to the existence of a minimal length. In fact, a minimal length actually appears in almost all proposed theories of quantum gravity. For this reason theories formulated in a minimal-length scenario are considered to be effective theories of quantum gravity [17, 18, 19, 20].
@@
In 1999, Yoon-Ho Kim and Yanhua Shih conducted an experiment whose results apparently suggested a violation of the HUP [21]. However, G. Rigolin pointed out that there is, in fact, no violation of the HUP, because the HUP is derived for particles that are non-correlated (non-entangled). In the Kim and Shih’s experiment, the photons of the pair are correlated (entangled) when one of the physical slits is replaced by a virtual slit [^2] The interaction of the photon with a physical slit destroys the correlation between the photons of the pair.. This indicates that the canonical HUP is no longer directly applicable due to the modification caused by quantum entanglement.[22, 23]
@@
An immediate question arising from the previous considerations is how the quantum entanglement modifies a generalized uncertainty principle (GUP), in others words, what is the effect caused by quantum entanglement at the minimal-length scale. The answer for that question is important in order to know the role of the quantum entanglement at the minimal-length scale (maybe appearing at the Planck scale) or in the early Universe. Unfortunately, this issue has been little considered in the literature. G. Blado, F. Herrera and J. Erwin [24] have studied the inseparability conditions with the most usual GUP correction, whereas D. Park[25] has used a coupled harmonic oscillator in order to find the effects of the quantum entanglement with a linear GUP.
@@
The outline of this paper is as follows. In Section 2 we obtain an expression of the uncertainty principle for entangled states which is independent of the chosen (generalized) uncertainty principle. In Section 3 we find the modified uncertainty principle for a pair of identical particles regarding the main proposals of GUP’s: Kempf, Mangano and Mann GUP (KMM-GUP), Ali, Das and Vagenas GUP (ADV-GUP), Pedram GUP and Nouicer GUP. In Section 4 we estimate an upper bound for the minimal-length value. We present our conclusions in the Section 5.
@@ 2 Uncertainty principle for entangled states
The position and momentum linear operators of the i-th particle, which act on the state vectors $\lvert\psi\rangle\in\cal{E}$, are the extensions $\tilde{Q}_{i}$ and $\tilde{P}_{i}$ defined as
@@
The relation (6) is general. It does not depend on whether the system of particle is entangled or not.
@@
Next, we briefly review the Rigolin’s result for a two particles system.
Initially, we define the functions
which are called quantum covariance functions (QCF). By definition, QCF’s vanish if and only if the system is separable [28]. Therefore (7) and (8) are zero for any not entangled quantum system. From the definitions of $\Delta Q$ and $\Delta P$ we have [^3] Note that $[\tilde{Q}_{1},\tilde{Q}_{2}]=0$ and $[\tilde{P}_{1},\tilde{P}_{2}]=0$.
@@
From now on, we omit the subscript $\psi$ for the sake of simplicity, whenever this does not cause any confusion. Thus,
@@
Using the QCF’s, Eqs. (7) and (8), we have
@@ 3 Uncertainty principle for entangled states in different minimal-length scenarios
In this section we consider a system of two entangled identical particles whose momenta have the same value but opposite directions, that is, $\vec{p}_{1}=-\vec{p}_{2}$, just as in the Kim and Shih’s experiment [21]. Therefore, in this case [^4] Note that $p_{1}=\lvert\vec{p}_{1}\rvert$ and $p_{2}=\lvert\vec{p}_{2}\rvert$, therefore $p_{1}=p_{2}$. $\left\langle\hat{p}_{1}\right\rangle=\left\langle\hat{p}_{2}\right\rangle$. Moreover, such a consideration also allows us to estimate, in the next section, an upper bound for the value of the minimal length based on the experimental results obtained by Kim and Shih.
@@ 3.1 Heisenberg uncertainty principle
Hence [^5] Note that $[\tilde{Q}_{i},\tilde{P}_{j}]=0$, for $i\neq j$.
@@
Substituting Eq. (23) into Eq. (20), and assuming [^6] We are going to assume the same for the next GUP’s. $\left(\Delta Q_{1}\right)^{2}=\left(\Delta Q_{2}\right)^{2}$ and $\left(\Delta P_{1}\right)^{2}=\left(\Delta P_{2}\right)^{2}$, we obtain
$\Delta Q_{i}\Delta P_{i}\geq\frac{\hbar}{4},$ (24)
where $i=1,2$. The result (24) shows that for a system of two entangled identical particles the HUP is modified. Such an outcome is not new, it was already obtained by G. Rigolin in 2002 [22] and then in 2016 [23].
@@ 3.2 KMM GUP
$[\tilde{Q},\tilde{P}]=i\hbar\left[2+\beta\left(\tilde{P}_{1}^{2}+\tilde{P}_{2}^{2}\right)\right]=2i\hbar\left(1+\beta\tilde{P}_{i}^{2}\right).$ (27)
@@ 3.3 ADV GUP
$[\tilde{Q},\tilde{P}]=i\hbar\left[2-2\alpha\left(\tilde{P}_{1}+\tilde{P}_{2}\right)+4\alpha^{2}\left(\tilde{P}_{1}^{2}+\tilde{P}_{2}^{2}\right)\right].$ (31)
@@
$\langle[\tilde{Q},\tilde{P}]\rangle=i\hbar\left[2-2\alpha\left(\left\langle\tilde{P}_{1}\right\rangle+\left\langle\tilde{P}_{2}\right\rangle\right)+4\alpha^{2}\left(\left\langle\tilde{P}_{1}^{2}\right\rangle+\left\langle\tilde{P}_{2}^{2}\right\rangle\right)\right],$ (32)
$\langle[\tilde{Q},\tilde{P}]\rangle=2i\hbar\left[1-2\alpha\left\langle\tilde{P}_{i}\right\rangle+4\alpha^{2}\left\langle\tilde{P}_{i}^{2}\right\rangle\right].$ (33)
@@
where $\gamma\prime:=-2\alpha\left\langle\tilde{P}_{i}\right\rangle+4\alpha^{2}\left\langle\tilde{P}_{i}\right\rangle^{2}$.
@@
It is important to note that the maximal uncertainty in the momentum is not changed, $\left(\Delta P_{i}\right)_{max}=\frac{1}{2\alpha}$. But this was to be expected because $\left(\Delta P_{i}\right)_{max}$ does not depend on $\hbar$.
@@ 3.4 Pedram GUP
With the intention of remedying some problems arising from KMM-GUP and ADV-GUP - such as the absence of a maximal momentum required in theories of doubly special relativity (DSR) and of commutative geometry, and also the lack of a valid approach for all order in the parameter related to the minimal length - P. Pedram has proposed a GUP [31, 32] based on the commutation relation given by
@@
$\Delta Q_{i}\Delta P_{i}\geq\frac{\hbar}{4}\frac{1}{\left[1-\beta\left(\Delta P_{i}\right)^{2}-\gamma\right]},$ (39)
where again $\gamma:=\beta\left\langle\tilde{P}_{i}\right\rangle^{2}$.
@@
3.5 Nouicer GUP
In the canonical field theory, in the context of non-commutative coherent states representation and field theory on non-anticommutative superspace the Feynman propagator displays an ultra-violet (UV) cut-off of the form $e^{-\beta p^{2}}$ [34, 35, 36, 37]. In consequence, K. Nouicer has proposed an exponential all orders GUP [38, 39] based on the commutation relation given by
@@ 3.5 Nouicer GUP
$\Delta Q_{i}\Delta P_{i}\geq\frac{\hbar}{4}e^{\beta\left(\Delta P_{i}\right)^{2}+\gamma},$ (44)
where once again $\gamma:=\beta\left\langle\tilde{P}_{i}\right\rangle^{2}$.
Therefore, the modified Nouicer-GUP (44) introduces a minimal uncertainty given by
@@
Finally, we note that $\left(\Delta Q_{i}\right)_{min}$ is also twice smaller than for a non-entangled system of a pair of identical particles.
@@ 3.6 Minimal length
If a non-zero minimal uncertainty in position can be interpreted as a minimal length then the previous results show that the minimal length for two entangled identical particles can be twice smaller than for a separable system. It is clear this statement must be taken with care because a minimal length should be a constant, that is, it must not depend on the physical system, it is a quantum gravitation effect. In fact, the minimal length should be an invariant as well as the light speed is. The answer for that apparent contradiction is possibly because the system is made up of two particles. C. Quesne and V. M. Tkachuck[40] have claimed that for a system composed by $N$ particles the effective parameter related to the minimal length, $\beta$, is reduced by a factor $\frac{1}{N^{2}}$. Hence, $\frac{\hbar}{2}\sqrt{\beta}$ is not the correct minimal uncertainty in position $\left(\Delta Q_{i}\right)_{min}$ for the modified KMM-GUP, but $\frac{\hbar}{2}\sqrt{\beta_{i}}$, where [40, 41]
@@
The reader might want to claim that the minimal length is actually described by the minimal uncertainty of the entangled system. However, according to G. Rigolin[23] we can presume that the minimal uncertainty for a system of $N$ entangled identical particles is reduced by $\frac{1}{N}$. Since, in principle, there is no limit for the number of particles for a entangled system, there would also be no limit for the minimal length.
@@ 4 Upper bound for the minimal-length value
$0.16\Delta P^{s}_{2}=\frac{\hbar}{2}\left[1+\beta\left(\Delta P^{s}_{i}\right)^{2}+\gamma\right].$ (48)
@@
$l_{min}\leq\hbar\sqrt{\frac{3.58\times 10^{-2}}{\hbar^{2}\eta}}\leq\hbar\sqrt{\frac{3.58\times 10^{-2}}{\hbar^{2}}}\approx 1.9\times 10^{-4}\textit{m}.$ (52)
@@ 5 Conclusion
In this work we find the non-zero minimal uncertainties induced by the main proposals of GUP’s (KMM, ADV, Pedram and Nouicer) which are modified due to the quantum entanglement of a system of two identical particles. In principle, our results have pointed out that the minimal uncertainties are reduced by half for a system of two entangled identical particles independently of the GUP. Hence, if a non-zero minimal uncertainty in position can be interpreted as a minimal length then the quantum entanglement reduces by half the minimal length. However, the minimal length must not depend on the physical system. We overcome this apparent paradox by using the Quesne and Tkachuck’s proposal for a system composed. Consequently, despite the quantum entanglement to change the GUP, the minimal length does not change. Based on our results and using the reference [23] we can expect that the apparent minimal uncertainty for a system of $N$ entangled identical particles is reduced by $\frac{1}{N}$, nonetheless the minimal length does not change because the effective parameter $\beta$ is also reduced by a factor $\frac{1}{N^{2}}$.